Keep the 15 kg of stack at the 2.00-meter mark on the left and move the...

  1. Keep the 15 kg of stack at the 2.00-meter mark on the left and move the 20 kg stack to a position of 1.25 m from the fulcrum to the right. Predict the position on the right side of the fulcrum where you must place a 5 kg brick to return to equilibrium. Show your work and record your answer in the box provided in the quiz. Check your prediction in the simulation.
Answer #1

let

m1 = 15 kg, d1 = 2 m
m2 = 20 kg, d2 = 1.25 m
m3 = 5 kg , d3 = ?

As the system is in equilibrium net torque and net force acting on the system must be zero.

Apply net torque about fulcrum = 0

m1*g*d1*sin(90) - m2*g*d2*sin(90) - m3*g*d3*sin(90) = 0

m1*d1 - m2*d2 - m3*d3 = 0

d3 = (m1*d1 - m2*d2)/m3

= (15*2 - 20*1.25)/5

= 1.00 m <<<<<<<<--------------------Answer

Note: counter clockwise torque is positive and clockwise torque is negative.